During a laboratory experiment, a 3.81-gram sample of NaHCO3 was thermally decomposed. In this experiment, carbon dioxide and water vapors escape and are combined to form carbonic acid. After decomposition, the sample weighed 2.86 grams. Calculate the percentage yield of carbonic acid for the reaction. Describe the calculation process in detail.

NaHCO3 → Na2CO3 + H2CO3
Answer #1

molecular weight of NaHCO3 = 84 g/mol (23 + 1 + 12 + 3 × 16)

molecular weight of H2CO3 = 62 g/mol (1 × 2 + 12 + 3 × 16)

3.81 gm NaHCO3 = 3.81 g / 84 g.mol-1= 0.045 mol

2.86 gm NaHCO3 = 2.86 g / 84 g.mol-1= 0.034 mol

Moles of NaHCO3 reacted = 0.011 mol

Balanced equation: NaHCO3 = 0.5 Na2CO3 + 0.5 H2CO3​​​​​​

1 mol of NaHCO3 produces 0.5 mol H2CO3​​​​​​

Moles of H2CO3 produced = 0.5 × 0.011 = 0.0055 mol = 0.0055 mol × 62 g/mol = 0.341 g

Actual yield = 0.341 g

If whole of NaHCO3 have reacted then H2CO3 produced = 0.045 mol × 0.5 × 62 g/mol = 1.395 g

Theoretical yield = 1.395 g

% yield = actual yield / theoretical yield × 100 = 0.341 / 1.395 × 100 = 24.44%